OpenStudy (anonymous):

does the integral of 1/x+ srqtx from 1 to infinity converge?

OpenStudy (anonymous):

i doubt it

OpenStudy (anonymous):

the degree of the denominator has to be larger than the degree of the numerator by more than 1

OpenStudy (anonymous):

is it \[\int_1^\infty\left({1\over x}+\sqrt{x}\right)dx\] OR \[\int_1^\infty{dx\over x+\sqrt{x}} \]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

@satellite73 aha... @peggiepenguin next time please use the "equation editor on this website" or "TEX" to enter your equations.. so we know what you are talking about!!

OpenStudy (anonymous):

so, it DOES converge coz, the denominator increases continuously and hence, the function decreases continuously

OpenStudy (anonymous):

oh no!!

OpenStudy (anonymous):

oh yes! :D

OpenStudy (anonymous):

just because the function decreases does not mean the integral converges that is a necessary but not a sufficient condition it must decrease fast enough

OpenStudy (anonymous):

simple example \[\int_1^{\infty}\frac{dx}{x}\] does not converge

OpenStudy (anonymous):

neither does the one above

OpenStudy (anonymous):

aah.. jumping to conclusions too soon. \[ u=\sqrt{x}\implies u^2=x\implies 2udu=dx\\ \int_1^\infty\frac{2u}{u^2+u}du\\ =2\int_1^\infty {du\over u+1}=2[\ln (u+1)]_1^\infty=\boxed{\infty} \]